HELLO: A 35% acid solution needs to be mixed with a 70% solution to create a 20 liter mixture that is 40% acidic. How many of each solution is required?
Howard Wilcox
howdywilcox@comcast.net

Answer by checkley79(3341)

.70x+.35(20-x)=.40*20
.70x+7-.35x=8
.35x=8-7
.35x=1
x=1/.35
x=2.857 liters of the 70% solution is used.
20-2.857=17.143 liters of 35% solution is used.
Proof:
.70*2.857+.35*17.143=.40*20
2+6=8
8=8